A stock had returns of 8%, 14%, and 2% for the past three years. Assuming these annual returns are normally distributed, what is the probability that this stock will earn at least 20% in any one given year?

a. 0.5%

b. 1%

c.1.5%

d.2.5%

e.5%
Answer #1

Average return = (8% + 14% + 2%) / 3 = 8%

Total squared deviation = (0.08 - 0.08)^2 + (0.14 - 0.08)^2 + (0.02 - 0.08)^2
= 0 + 0.0036 + 0.0036
= 0.0072

Standard deviation = (0.0072 / (3 - 1))^(1/2) = 0.06 or 6%

Upper point of 95% probability range = 8% + [2 × 6%] = 8% + 12% = 20%

Probability associated with upper tail of 95% range = ½ × (100% - 95%)
= ½ × 5%
= 2.50%

Probability that this stock will earn at least 20% in any one given year = 2.5%